OpenStudy (anonymous):

please help me ! A number being successively divided by 3, 5 and 8 leaves remainders 1, 4 and 7 respectively. Find the remainders if the order of divisors be reversed explain it with correct procedure.

ganeshie8 (ganeshie8):

say the starting number is \(x\)

ganeshie8 (ganeshie8):

then we have : \( x = 3i + 1 \\ i = 5j + 4 \\ j = 8k + 7 \)

ganeshie8 (ganeshie8):

eliminating \(j\) from first two equations gives : \(x = 3(5j+4) + 1 = 15j + 13\)

ganeshie8 (ganeshie8):

eliminating \(j\) gives : \(x = 15(8k+7) + 13 = 120k + 118\)

ganeshie8 (ganeshie8):

\(x = 120k + 118\) divide this in reverse order and see what u get